# Covariance-Based Device Activity Detection with Massive MIMO for Near-Field Correlated Channels

### v1

@@ V-B Numerical Validation of Detection Performance Analysis
In this subsection, we verify the detection performance analysis in Section IV. We consider the device activity detection problem under Case I and Case II and perform MLE model (10) in both cases. In order to demonstrate the impact of channel correlation matrices on the detection performance as compared to the traditional uncorrelated Rayleigh fading channels, we set for all for a fair comparison.
@@
In Fig. 3, we show the detection performance for different values of with and The “probability of error” in the figure is defined as the point where the probability of missed detection (PM) and probability of false alarm (PF) are equal. We observe that increasing can improve the detection performance in both cases. When is small, the detection performance in Case I is better, and when is large, the detection performance in Case II is better. This is because is rank-deficient in Case I and full-rank in Case II, which makes the signal statistical dimension (defined in (42) and (43)) in Case II larger than that in Case I when is large. In other words, in Case I, some dimensions of the received signal are actually useless. We can also observe from Fig. 3 that in Case I, when the larger is, the better the detection performance is. This also verifies when is larger, the rank of is larger, and thus the signal statistical dimension is larger, and the detection performance is better.
@@
Fig. 4 plots the detection performance for various values of with parameters and It is observed that the detection performance of both cases deteriorates as increases. Furthermore, when is large, for instance, the detection performance in Case I is superior to that of Case II. This is due to when is large, the signal statistical dimensions are equal in both cases; however, in Case I, the true activity vector is more easily identifiable compared to Case II, as discussed in Section IV-B. Additionally, we also observe that in Case I, a smaller tends to result in better detection performance for larger This can be explained by the fact that when is smaller, the cosine similarity in (54) is likely to be lower, indicating a weaker correlation between different devices, which in turn leads to improved detection performance.
Fig. 5 shows the detection performance for various values of with and We observe that as increases, the detection performance in Case I with values of and significantly outperforms that of Case I with and Case II. This suggests that when is large and the rank of is large in Case I, the detection performance is superior compared to the case when the rank of is low in Case I, and it is also superior to Case II.
V-C Detection Performance Comparison with Benchmarks
@@
AMP [31] and OAMP [34]: Both algorithms are designed for joint activity detection and channel estimation under correlated Rayleigh fading channels. OAMP is an improved version of AMP, which is applicable to more types of signature sequences. In the scenario of Rician fading channels, i.e., we ignore the LoS channel and assume the prior distribution to be in both methods. Notice that both methods require the knowledge of the probability that a device is active, i.e., but the MLE does not require this prior information.
@@
VI Conclusion
This paper investigates the device activity detection problem for NFC, where the near-field channel can be modeled as correlated Rician fading. The corresponding device activity detection problem is formulated as an MLE problem. Two efficient algorithms for solving the MLE problem, i.e., exact and inexact CD, are proposed. In addition, an analysis of the detection performance of MLE is provided. Simulation results demonstrate the efficiency of the proposed algorithms and verify the correctness of the detection performance analysis.
@@ Appendix A Proof of Theorem 1
Furthermore, we have a quadratic upper bound on [57, Proposition A.24]:
@@
and the properties of the projection operator Similar techniques are used in the proof of [8, Proposition 4]. Substituting into (63) and combining it with (A), we obtain (29). This completes the proof of Theorem 1.

### v2

@@ VI-B Numerical Validation of Detection Performance Analysis
In this subsection, we verify the detection performance analysis in Section IV. We consider the device activity detection problem under Case I and Case II and perform MLE model (10) in both cases. Notice that the key difference between Case I and Case II lies in the channel correlation matrices. In order to highlight the impact of the channel correlation matrices on detection performance while eliminating the effect of channel means, we set the channel means to zero in this subsection. The termination condition of the inexact CD algorithm is set as
@@
In Fig. 4, we show the detection performance for different values of with and The “probability of error” in the figure is defined as the point where the probability of missed detection (PM) and probability of false alarm (PF) are equal. We observe that increasing can improve the detection performance in both cases. When is small, the detection performance in Case I is better, and when is large, the detection performance in Case II is better. This is because is rank-deficient in Case I and full-rank in Case II, which makes the signal statistical dimension (defined in (42) and (43)) in Case II larger than that in Case I when is large. In other words, in Case I, some dimensions of the received signal are actually useless. We can also observe from Fig. 4 that in Case I, when the larger is, the better the detection performance is. This also verifies when is larger, the rank of is larger, and thus the signal statistical dimension is larger, and the detection performance is better. The crossover for and is because when is small, the signal statistical dimension is the same for both cases, as it cannot exceed When is large, the larger the is, the larger the rank of will be, and consequently, the larger the signal statistical dimension. Therefore, the detection performance for is better than that for
@@
Fig. 5 plots the detection performance for various values of with parameters and It is observed that the detection performance of both cases deteriorates as increases. Furthermore, when is large, for instance, the detection performance in Case I is superior to that of Case II. This is due to when is large, the signal statistical dimensions are equal in both cases; however, in Case I, the true activity vector is more easily identifiable compared to Case II, as discussed in Section IV-B. Additionally, we also observe that in Case I, a smaller tends to result in better detection performance for larger This can be explained by the fact that when is smaller, the cosine similarity in (54) is likely to be lower, indicating a weaker interference between different devices, which in turn leads to improved detection performance.
Fig. 6 plots the detection performance for various values of with parameters and Notice that the sparsity level of active devices ranges from to We observe that the detection performance in both cases deteriorates as increases, exhibiting a similar trend to that as in Fig. 5 when increases. When is large, the detection performance in Case I is superior to that in Case II. Additionally, within Case I, a smaller value corresponds to better detection performance for large This is because, as discussed in Section IV-B, the interference between devices in Case I is less than that in Case II. Moreover, in Case I, a smaller value corresponds to a smaller cosine similarity in (54), resulting in less interference between devices.
Fig. 7 shows the detection performance for various values of with and We observe that as increases, the detection performance in Case I with values of and significantly outperforms that of Case I with and Case II. The detection performance of Case I with is worse than that of Case I with and when is large, because the signal statistical dimension in Case I with is lower than that in Case I with and The detection performance of Case II is worse than that of Case I with and because, in Case I, the identifiability of the true activity vector is easier to hold, and there is less mutual interference between different devices.
VI-C Detection Performance Comparison with Benchmarks
@@
AMP [32] and OAMP [35]: Both algorithms are designed for joint activity detection and channel estimation under correlated Rayleigh fading channels. OAMP is an improved version of AMP, which is applicable to more types of signature sequences. Notice that while AMP and OAMP can be applied under correlated Rayleigh fading channels, they cannot be directly applied to correlated Rician fading. The reason is that the denoiser and state evolution in these algorithms assume zero channel means, which cannot be directly extended to scenarios with nonzero channel means. In the scenario of Rician fading channels, i.e., we ignore the mean and assume the prior distribution to be in both methods.
Notice that both methods require the knowledge of the probability that a device is active, i.e., but the MLE does not require this prior information.
@@
The termination condition of the inexact CD algorithm is set as and we allow other benchmarks to run sufficiently long until they converge.
Notice that AMP [32] and OAMP [35] are CS-based methods, while the MLE proposed in this paper and mismatched MLE are covariance-based methods. The key difference between the CS-based method and the covariance-based method is that the former jointly estimates the channels and device activities, whereas the latter estimates only the device activities. Other covariance-based methods include SPICE [30], which can be applied to the Rayleigh fading scenario but encounters technical difficulties in the Rician fading scenario.
@@
VI-D Joint Device Activity and Data Detection
In this subsection, we study the joint device activity and data detection problem and demonstrate the detection performance of the MLE model (60).
Fig. 11 plots the detection performance for various values of with and We observe from Fig. 11 that when is small, the smaller the number of information bits the better the detection performance. In addition, as increases, for bit and the detection performance of the MLE model improves and gradually becomes equal. Note that when there is no data transmission, and only activity detection is performed. This indicates that in the joint device activity and data detection problem, transmitting bits of information (when is small) does not significantly deteriorate the detection performance of the MLE model, especially when is large.
VII Conclusion
This paper investigates the device activity detection problem for NFC, where the near-field channel can be modeled as correlated Rician fading. The corresponding device activity detection problem is formulated as an MLE problem. We present the exact CD algorithm and propose the inexact CD algorithm with improved numerical stability for solving the MLE problem. In addition, an analysis of the detection performance of MLE is provided. Simulation results demonstrate the efficiency of the presented algorithms and verify the correctness of the detection performance analysis.
@@ Appendix A Proof of Theorem 1
Furthermore, we have a quadratic upper bound on [60, Proposition A.24]:
@@
and the properties of the projection operator Similar techniques are used in the proof of [9, Proposition 4]. Substituting into (69) and combining it with (A), we obtain (29). This completes the proof of Theorem 1.